OpenStudy (anonymous):

4 greater than or equal to x/-2? HELP

OpenStudy (jdoe0001):

is ... it solve for "x" well... , just cross-multiply :) \(\bf 4\ge \cfrac{x}{-2}\\ \quad \\ \textit{let us multiply both sides by }\quad -2\\ -2(4) \ge -2\left(\cfrac{x}{-2}\right)\\ \quad \\ \textit{in an inequality, when multiplying/dividing/exponentializing}\\ \textit{by a negative number, you have to } \text{flip}\textit{ the sign}\\ \quad \\ -2(4) \boldsymbol{\color{red}{\ge}} -2\left(\cfrac{x}{-2}\right)\implies -8 \boldsymbol{\color{red}{\le}} x\)

OpenStudy (anonymous):

Thanks guys i figured it out and got x greater than or equal to negavtive 8